Translate the problem into a pair of linear equations in two variables. Solve the equations using either elimination or substitution. State your answer for the specified variable.
Walt made an extra $5000 last year from a part-time job. He invested part of the money at 6% and the rest at 10%. He made a total of $420 in interest. How much was invested at 10%?

Answer by stanbon(75887)

Translate the problem into a pair of linear equations in two variables. Solve the equations using either elimination or substitution. State your answer for the specified variable.
Walt made an extra $5000 last year from a part-time job. He invested part of the money at 6% and the rest at 10%. He made a total of $420 in interest. How much was invested at 10%?
-----------------
Quantity Equation: s + t = 5000
Interest Equation:0.06s + 0.10t = 420
---------------------------------------------
Modify the equations for elimination:
Multiply thru the 1st by 6 ; Multiply thru the 2nd by 100
6s + 6t = 6*5000
6s + 10t = 42000
-----------------------
Subtract the 1st equation from the 2nd and solve for "t"
4t = 12000
t = $3000 (amount invested at 10%)
==============================================
Cheers,
Stan H.
Reply to stanbon@comcast.net